OpenStudy (anonymous):

simple but confused me :( find the slope of f(x) at the point where it crosses the x-axis f(x)=1-e^x ------------ I'm a little bit confused, now the function or graph is going to cross the x-axis at the points of its roots, i.e when f(x)=0 is that right :d so 1-e^x=0 e^x=1 x=ln1 x=0 so the root is zero, and the graph is going to cross the x-axis when x=0 ??? the slope at that point is going to be the derivative of the function at the same point, f'=e^x so the slope : =e^0 = 1 ------- is my work OK, or there is a big misunderstanding :(

OpenStudy (anonymous):

here we go, the king is around :) how r u man :)

OpenStudy (anonymous):

who r u referring to? i hope not me because this is all a guess.

OpenStudy (anonymous):

kidding!!

OpenStudy (anonymous):

check your derivative... that's not right.

OpenStudy (anonymous):

is it -e^x ?

OpenStudy (anonymous):

yes. so the slope is -1

OpenStudy (anonymous):

what about the equation of the tangent line to the curve at this point !

OpenStudy (anonymous):

what are the x, y coordinates of the point in question?

OpenStudy (anonymous):

the question says when the graph cross the x-axis, there is no x,y coordinates

OpenStudy (anonymous):

every point on the graph has an x,y coordinate

OpenStudy (anonymous):

this is my confusing, how to get the x, y of the point, I'm suppose to get the coordinates by finding the point where the graph cross the x-axis

OpenStudy (anonymous):

I'm really confused :(

OpenStudy (anonymous):

ok... the point in question is the point where the graph crosses the x-axis.... you already have the x-coordinate of where that happens... what is it?

OpenStudy (anonymous):

y=0 x=0

OpenStudy (anonymous):

yes.....

OpenStudy (anonymous):

we have the slope -1

OpenStudy (anonymous):

so, you have a point (0, 0), and a slope at that point m=-1... how d u use this info to come up with the equation of a line?

OpenStudy (anonymous):

is it y=mx+b so 0=-1(0)+b lool

OpenStudy (anonymous):

horrible work I now :D

OpenStudy (anonymous):

do you know point-slope form for the equation of a line? y -y0 = m*(x - x0)

OpenStudy (anonymous):

you don't have to use y=mx + b..

OpenStudy (anonymous):

y -y0 = m*(x - x0) where (x0, y0) is the point you know and m is the slope you know...

OpenStudy (anonymous):

y-0=-1(x-0)

OpenStudy (anonymous):

ok... you can leave it like that. but do you think you can simplify that a bit?

OpenStudy (anonymous):

yes I'll try now

OpenStudy (anonymous):

y-0=-x-0

OpenStudy (anonymous):

how about y = -x

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

other than your derivative being miscalculated your work looks fine... good work man...

OpenStudy (anonymous):

you r just great man, the way you teach is amazing, wish all the best in ur life :)

OpenStudy (anonymous):

and I hope me guessing how to do this by the seat of my pants yielded correct results.

OpenStudy (anonymous):

kidding!!!

OpenStudy (anonymous):

lool

OpenStudy (anonymous):

thanx for ur time

OpenStudy (anonymous):

np & yw...

OpenStudy (anonymous):

\[f(x) =1-e^x\\ f(0) = 0\\ \text { You need to find } f"(0).\\ f'(x)= - e^x\\ f'(0) = -1 \]

OpenStudy (anonymous):

You need to find f'(0) There was a misprint above.